OpenStudy (dtan5457):

Someone check my work? Product to get trinomial? (1/2y^2-1/3y) and (12y+3/5)

OpenStudy (dtan5457):

I got 6y^3-(37/10y^2)-1/5y

OpenStudy (anonymous):

\(\dfrac{1}{2y^2}-\dfrac{1}{3y}\) and \(12y +\dfrac{3}{5}\) right?

OpenStudy (dtan5457):

the y's are for the whole fraction but yea thats right

OpenStudy (anonymous):

can you use draw box below to draw the problem out? like

OpenStudy (anonymous):

then you are correct. :)

OpenStudy (dtan5457):

thanks.